Electrostatics

n small drops of same size are charged to V volts each. if they coalesce to form a signal large drop, then its potential will be?

sagar singh
13 Years agoGrade

1 Answer

Sher Mohammad
12 Years ago
volume for both cases will be same so 4/3*pi*r^3*n=4.3*pi*R^3
=>R=n^(1/3)*r

total charge afterward is nQ
so potential k*nQ/(n^(1/3)*r)=n^(2/3)*kQ/r=n^(2/3)*V

Sher Mohammad
Askiitians Faculty